# FlashDMoE: Fast Distributed MoE in a Single Kernel

### v2

@@ Abstract
The computational sparsity of Mixture-of-Experts (MoE) models enables sub-linear growth in compute cost as model size increases, thus offering a scalable path to training massive neural networks. However, existing implementations suffer from low GPU utilization, significant latency overhead, and a fundamental inability to leverage task locality, primarily due to CPU-managed scheduling, host-initiated communication, and frequent kernel launches. To overcome these limitations, we develop FlashDMoE, a fully GPU-resident MoE operator that fuses expert computation and inter-GPU communication into a single persistent GPU kernel. FlashDMoE enables fine-grained pipelining of dispatch, compute, and combine phases, eliminating launch overheads and reducing idle gaps. Unlike existing work, FlashDMoE obviates bulk-synchronous collectives for one-sided, device-initiated, inter-GPU (R)DMA transfers, thus unlocking payload efficiency, where we eliminate bloated or redundant network payloads in sparsely activated layers. When evaluated on an 8-H100 GPU node with MoE models having up to 128 experts and 16K token sequences, FlashDMoE achieves up to 9$\times$ higher GPU utilization, 6$\times$ lower latency, 5.7$\times$ higher throughput, and 4$\times$ better overlap efficiency compared to state-of-the-art baselines—despite using FP32 while baselines use FP16. FlashDMoE shows that principled GPU kernel-hardware co-design is key to unlocking the performance ceiling of large-scale distributed ML.
@@
Figure 1: FlashDMoE performance.
@@ 1 Introduction
State-of-the-art large language models (LLMs), including DeepSeek-v3 [1], LLama4 [2], DBRX [3] and Snowflake Arctic [4], have adopted the Mixture-of-Experts (MoE) architecture for its computational efficiency and strong performance across many tasks. The traditional Transformer block consists of a self-attention module followed by a dense feed-forward network (FFN) [5]. In contrast, MoE architectures replace this single FFN with identically sized FFNs, otherwise known as experts, (Figure 2(b)). A trainable neural network, known as a gate function, sparsely activates these experts by dynamically routing input tokens to selected experts at runtime. This increase in model parameters (more FFNs) improves model quality without a corresponding increase in computational cost.
@@
Communication overheads in MoE. As MoE model sizes grow, GPU memory constraints prevent hosting all experts on a single device. The standard practice is to distribute experts across multiple GPUs using expert parallelism (EP), which requires token routing via many-to-many communication in AllToAll or AllGather [1, 4, 3, 6]. Another round of said many-to-many communication is also necessary for restoring the permuted tokens processed by experts to their original order within the sequence. Existing work has observed these communication operations taking 68% of total runtime [7, 8], during which the GPU is completely idle, unless the implementation explicitly overlaps with computation. This form of pipelining is challenging to achieve efficiently because it requires asynchronous GPU-driven communication and kernel fusion to maximize the overlap efficiency. Typically, inter-GPU communication APIs available in frameworks like PyTorch are not of this kind but instead are CPU-driven [9].
@@
Kernel launch overheads in MoE. The efficacy of communication overlap is further limited by the overhead of launching many kernels from the CPU. Specifically, existing implementations [10, 11, 12, 13] require launching a large number of kernels per a single layer pass (see Table 1). Frequent kernel launches negatively affect performance by: (1) creating non-deterministic kernel start times across GPUs, exacerbating straggler issues; (2) introducing unnecessary synchronization points, causing GPUs to wait on peers or the CPU before proceeding; and (3) incurring repeated global memory round trips at kernel boundaries. Although CUDA graphs [14] can partially mitigate the first issue in static workloads, they are incompatible with MoE’s dynamic expert routing patterns. Addressing the remaining issues requires novel solutions, which we provide in this work through complete kernel fusion and asynchronous device-initiated communication.
@@
Figure 3: Comparing FlashDMoE with state-of-the-art techniques that either do not overlap communication and computation (left, top) or do some overlap (left, middle). FlashDMoE is a persistent kernel that fuses all computation and communication of the MoE operator (left, bottom). FlashDMoE implements device-initiated computation (gate, expert FFN, scale) and communication tasks (right).
Table 1: Kernel Fusion Comparison. Our method is the first to fully fuse the DMoE layer into a single GPU kernel. We report GPU operations from profiling with Nsight Systems. We count within a single layer (Gate $\rightarrow$ Dispatch $\rightarrow$ Expert $\rightarrow$ Combine) on 2 A100s, where each GPU has 32 experts.
Works Launched GPU Ops
FlashDMoE 1
COMET [11] 33
Megatron-LM CUTLASS [12, 15] 85
Megatron-LM TE [12, 15] 261
Megatron-LM + DeepEP [1] 432
DeepSpeedMoE [10] 550
1.1 Our Contributions: DMoE in a single kernel
@@
To overcome these fundamental inefficiencies in state-of-the-art MoE models, we develop FlashDMoE, where we integrate all DMoE computation and communication tasks into a single persistent GPU kernel i.e., a kernel that remains active for the entirety of the MoE operator (Figure 3 bottom left). Instead of multiple kernel launches coordinated by the CPU, FlashDMoE requires launching only one kernel, significantly reducing the involvement of the CPU. Within the fused kernel, FlashDMoE implements a reactive programming model to achieve fine-grained parallelism and loosely coupled, non-blocking execution among tens of thousands of GPU threads.
@@
In-kernel Block scheduling and Tile parallelism. FlashDMoE implements tile-level parallelism, meaning it partitions input token matrices into smaller, independent units called tiles, which are processed by blocks but managed (scheduled and constructed) by warps. We specialize every thread block, except one, as processors to perform compute. In addition, we designate a dedicated Operating System (OS) block (4 warps) to perform administrative tasks of (1) scheduling computational work to processors (scheduler), and (2) decoding computational tasks from messages received from other GPUs (subscriber). This design allows FlashDMoE to dynamically assign tasks to GPU blocks based on readiness, ensuring that no GPU SM remains idle throughout the lifetime of the DMoE operator. FlashDMoE selects tile dimensions to maximize GPU arithmetic intensity while still benefitting from a high-degree of parallelism.
@@
Asynchronous and payload-efficient communication. By redesigning the MoE operator from the ground up, FlashDMoE resolves fundamental inefficiencies inherent in the conventional MoE execution pipeline. One notable inefficiency is token padding during communication. To simplify programming complexity and due to symmetry constraints of collective communication APIs, existing implementations have to zero-pad token payloads to match predefined buffer sizes. This occurs when tokens are asymmetrically routed to experts, resulting in GPUs receiving much less than the expected capacity. However, these null payloads waste communication bandwidth, bloat data transfer latency and may lead to unnecessary computations on null matrices in some implementations. FlashDMoE introduces payload-efficient communication by sending non-padded tokens only to GPUs with actively selected experts, conserving both communication and computational resources.
@@
Technical challenges. Realizing the single-kernel design of FlashDMoE required solving several technical challenges to achieve high performance: (1) lightweight computational dependency management; (2) navigating optimal SM occupancy configurations; (3) implementing in-device BLAS operations; (4) minimizing inter- and intra-device synchronization overheads; (5) implementing transfer-awareness by leveraging DMA over Unified Virtual Addressing (UVA) when available. In addressing these challenges, FlashDMoE’s design presents a radical departure from traditional synchronous AllToAll collectives, where GPUs exhibit significant idle time during layer execution. For device-initiated communication, FlashDMoE uses NVSHMEM [16] to establish a global address space across all GPUs to achieve the aforementioned Direct Memory Access (DMA) or Remote DMA (RDMA) communication. For in-device BLAS, FlashDMoE develops custom high-performance GEMM operations via CUTLASS [17].
Results. We evaluate FlashDMoE across multiple GPUs split across multiple nodes. Our evaluations show that FlashDMoE achieves 6$\times$ latency speedup, 9$\times$ higher GPU utilization, 4$\times$ better weak scaling efficiency and 5.7$\times$ increased throughput compared to state-of-the-art implementations. We project these performance gains becoming even better in multi-node scenarios, where inter-node communication occurs using lower bandwidth inter-node links (e.g., RDMA, Infiniband).
@@ 2 Motivation
2.1 Synchronous Communication and Stragglers
Figure 4: Overlapped Schedule (bottom) showing how idle time from the sequential schedule (top) is repurposed for computation. FlashDMoE implements the overlapped schedule.
AlltoAll or AllGather communication as currently used in MoE frameworks is a synchronous collective operation, whose completion requires the participation of all involved GPUs. Here, disparities in processing speeds or kernel scheduling among GPUs induce a straggler effect detrimental (Figure 4) to (1) the collective operation’s performance and (2) E2E performance, as stalled GPUs cannot proceed to downstream dependent or independent tasks until the collective terminates. Specifically, as shown in Figure 15, for distributed training of a 1.3B GPT-3 MoE model across 32 A100 GPUs, we see P95 communication performance degradation of 1.32X when compared to the mean actual kernel time from Figure 15(b). This performance reduction is rather tame as the underlying hardware is a supercomputer well-tuned against “software jitter” [18]. However, we observe a more severe p95 performance loss of 11X in a single-node Virtual Machine (VM). In line with prior HPC works [19, 20], we argue that obviating the inherent barrier in this synchronous collective communication would allow GPUs to repurpose this observed idle time for downstream computation as depicted in Figure 4.
Table 2: Straggler Delay within Synchronous All-to-All communication. We capture the distribution of delay induced by stragglers across many steps. Let Actual Time $t_{a}$ denote the fastest kernel execution time across all GPUs, and Total Time $t$ be the maximum recorded step time. We define $Delay$ as the maximum difference between $t$ and $t_{a}$. Note $Delay$ is idle time. For the 1x8 V100, we profile 1750 steps and 600 steps for the 8x4 A100. See Figure 15 for the raw distribution.
System # Nodes # GPUs Median p95
Commercial VM (V100) 1 8 3.1x 11.4x
Supercomputer (A100) 8 32 1.09x 1.32x
2.2 Kernel launch overhead.
@@ 2.1 Synchronous Communication and Stragglers
Figure 5: 5(a) shows GPU utilization averaged across 100 MoE forward passes on 2 NVLinked A100s with 300 GB/s unidrectional bandwidth. Despite the high-bandwidth interconnect, we observe up to 90% idle time, which we attribute to kernel launch gaps and non-overlapping communication.
We compare the kernel launch overheads between FlashDMoE and existing baselines. Table 1 shows the number of kernel launches during a single forward pass: FlashDMoE launches exactly one persistent kernel, while the baselines launch up to 550 short-lived kernels to perform the same computation. Figure 5 provides a visual comparison using CUDA API traces captured by NSight Systems, illustrating the difference between FlashDMoE and DeepEP. DeepEP exhibits numerous small CUDA API calls, with frequent stalls between individual operators, leading to increased GPU idle time (Figure 5(a)). In contrast, FlashDMoE maintains high GPU utilization by avoiding launch overhead and synchronization gaps—achieving 93.17% GPU utilization compared to 14% for DeepEP. See §4 for experimental results and §A for a discussion of related work.
@@ 3 Fused MoE Kernel Design
Figure 6: FlashDMoE Fused Kernel
Modern distributed MoE systems suffer from two limitations: (1) frequent many-to-many (AlltoAll or AllGather) collectives on the critical path, and (2) significant overhead from repeated kernel launches. We address these in FlashDMoE, a fully fused MoE operator implemented as a single persistent GPU kernel. Unlike previous approaches [11, 1, 10, 12, 21, 8, 22, 23, 24, 25, 26], FlashDMoE is the first solution to implement a completely fused Distributed MoE kernel, eliminating kernel launch overhead entirely by requiring only a single kernel launch (see Table 1).
@@ 4             $\mathbf{Dispatch}(T_{\phi},A)$
Algorithm 1  FlashDMoE Distributed MoE Fused Kernel
@@
Figure 7: DMoE Functional Dependencies Expressed as a Chain of Actor Interactions. We denote $S_{b}$, $S_{h}$, and $P$ as the Subscriber, Scheduler and Processor actors, respectively. For any actor $a\in\{S_{b},\>S_{b},\>P\}$, $a^{i}$ identifies an actor on GPU $i$. We define $D^{j}_{i}$ as the operator, where GPU $j$ dispatches packets of tiles to GPU $i$, This diagram expresses task dependencies at the granularity of a tile, namely $GEMM_{0}$, $GEMM_{1}$, combine and communication produce an output tile. Notifications occur as signals propagated through shared memory (subscriber $\leftrightarrow$ scheduler) or global memory (scheduler $\leftrightarrow$ processor or inter-GPU communication). Note one-sided inter-GPU transfers (packet or single tile) are coupled with a signal to notify $S_{b}^{j}$ on the receiving GPU $j$ of the message’s delivery.
@@
Actor-based model. The design of FlashDMoE is based on the actor model of concurrent computation [27, 28, 29]. We implement this model by specializing GPU thread blocks and warps into three distinct actor roles: (1) ProcessorE.1), (2) SubscriberE.3), and (3) SchedulerE.2). The Processor performs compute (GEMMs and element-wise operations) and tile communication. We use CUTLASS [17] as the underlying infrastructure for high-performance BLAS routines and NVSHMEM for kernel-initiated communication [16]. The Subscriber and Scheduler perform administrative functions. Specifically, the Scheduler assigns computational tasks to available thread blocks. Our key innovation is making the Scheduler both multithreaded, enabling high scheduling throughput, and work-conserving, ensuring consistently high GPU SM utilization. On the other hand, the Subscriber decodes tile packets from peer GPUs to task descriptors (§3.1). Of the $N$ thread blocks on a GPU, we specialize $N-1$ to adopt the Processor role. We specialize the last block as the Operating System (OS). Within this block, we specialize three warps for the Subscriber role and one warp for the Scheduler role. This split of thread blocks across actors is intentional: our goal is to use few resources for administrative tasks while reserving bulk of the resources for performing MoE computation tasks. Figure 6 summarizes the FlashDMoE architecture and its constituent actors, while Algorithm 1 gives a very close translation of the system in code. Note that $A\in\mathbb{R}^{S\times H}$ is the input token matrix; $O\in\mathbb{R}^{S\times H}$ the output matrix; and $X\in\mathbb{R}^{E\times H\times D}$ is a 3-D tensor of expert weights, where $E$ denotes the number of local experts for the executing GPU, $H$ is the embedding dimension, $D$ is the FFN intermediate dimension and $S$ is the sequence length. $T_{\phi}\in\left(\mathbb{R}^{2}\right)^{E\times C}$ is a routing table data structure, where $T_{\phi}\left(e,c\right)=(i,w)$ indicates that token $i$ at slot $c$ dispatches to expert $e$. $w$ is the combine weight (Equation 2) and $C$ is expert capacity. The tuple structure of $T_{\phi}$ is an implementation detail. $G_{\phi}\in\mathbb{R}^{S\times E}$ captures the affinity scores produced by the gate (Equation 3).
Figure 8: GPU Memory Hierarchy. The inverted pyramid (left) shows the load/store access latency [30, 31, 32]. The table above outlines the capacity for different memory tiers (for A100 GPUs). The shared memory and register capacity are static configurations for FlashDMoE. The right figure shows accessibility scopes: on-chip registers are scoped to a thread; on-chip shared memory is visible to all threads in a block; and off-chip global memory is accessible by all threads on device.
Inter-actor interactions in FlashDMoE. FlashDMoE decomposes MoE computation and communication at the granularity of a tile, a statically sized partition of a tensor, to achieve parallel execution and efficient overlap of tasks. Each tile maps to a discrete unit of work encapsulated by a task descriptor. The Subscriber decodes these task descriptors from the remote tile packets it receives. Concurrently, the Scheduler receives notifications about available tasks and dispatches them for execution to Processor actors that perform computations defined by these tasks, namely the feed-forward network (FFN) and expert-combine operations. Figure 7 show the chain of actor interactions, demonstrating how FlashDMoE enforces DMoE functional dependencies.
@@
Determining tile dimensions in FlashDMoE. Selecting appropriate tile dimensions in FlashDMoE is crucial to ensure efficient GPU utilization. An undersized tile underutilizes the GPU, while excessively large tiles create register pressure, causing performance-degrading register spills to local memory. After careful parameter sweeps, we choose tile dimensions of (128, 64). Our key insights are: increasing tile width significantly raises the register usage per thread, potentially triggering costly spills; increasing tile height without adjusting thread count increases workload per thread, harming performance. Raising the thread count per block beyond our fixed value of 128 threads reduces the number of concurrent blocks, negatively affecting SM occupancy. Larger thread-block sizes also increase overhead from intra-block synchronization (__syncthreads() barriers), further degrading performance. Thus, our chosen tile dimensions balance register usage, shared-memory constraints, and GPU occupancy to deliver optimal performance.

### v3

# FlashMoE: Fast Distributed MoE in a Single Kernel

@@ Abstract
[…]
@@
Figure 1: FlashMoE performance.
@@ 1 Introduction
State-of-the-art large language models (LLMs) [deepep, llama4, dbrx, arctic, openai2025gptoss] have adopted the Mixture-of-Experts (MoE) architecture for its computational efficiency and strong performance across a range of tasks. The traditional Transformer block consists of a self-attention module followed by a dense feed-forward network (FFN) [NIPS2017_3f5ee243]. In contrast, MoE architectures replace this single FFN (Figure 2(a)) with many identically sized FFNs, known as experts (Figure 2(b)). A trainable neural network, known as a gate function, sparsely activates these experts by dynamically routing input tokens to the experts selected at runtime. This increase in model parameters due to more FFNs improves model quality without the corresponding increase in computational cost.
@@
Communication overheads in MoE. As MoE model sizes grow, GPU memory constraints prevent hosting all experts on a single device. The standard practice is to distribute experts across multiple GPUs using expert parallelism (EP), which requires token routing via many-to-many communication primitives like AlltoAll [deepep, arctic, dbrx, 10.1145/3577193.3593704] (Figure 2(c)). Another round of AlltoAll communication restores the permuted tokens processed by experts to their original order in the sequence. AlltoAll communication is challenging to optimize on GPU networks and is highly sensitive to straggler delays — a phenomenon where a single straggler GPU delays all others from making progress [stragglar]. These communication operations can account for 68% of the total runtime [10.1145/3603269.3604869, MLSYS2024_339caf45], causing GPUs to remain idle (Figure 3, top left).
@@
Kernel launch overheads in MoE. To mitigate these communication bottlenecks, recent work pipelines computation with communication kernels (Figure 3, left middle). However, the effectiveness of these solutions is limited by the overhead of launching many kernels from the CPU. Specifically, existing implementations [pmlr-v162-rajbhandari22a, comet, megatron, fastermoe] launch a large number of kernels per a single layer pass (see Table 1). Frequent kernel launches negatively affect performance by: (1) creating non-deterministic kernel start times across GPUs, exacerbating straggler issues; (2) introducing unnecessary synchronization points, causing GPUs to wait on peers or the CPU before proceeding; and (3) incurring repeated global memory round trips at kernel boundaries. Although CUDA graphs [cuda_graphs_nvidia_blog] can partially mitigate the first issue in static workloads, they are incompatible with MoE’s dynamic expert routing patterns. Addressing the remaining issues requires novel solutions, which we provide in this work through complete kernel fusion and asynchronous device-initiated communication.
@@
Figure 3: Comparing FlashMoE with state-of-the-art techniques that either do not overlap communication and computation (left, top) or do some overlap (left, middle). FlashMoE is a persistent kernel that fuses all computation and communication of the MoE operator (left, bottom). FlashMoE implements device-initiated computation (gate, expert FFN, scale) and communication tasks (right).
@@
Our Contributions: distributed MoE in a single kernel. To overcome these fundamental inefficiencies in state-of-the-art MoE models, we develop FlashMoE a megakernel that integrates all MoE computation and communication tasks into a single persistent GPU kernel i.e., a kernel that remains active for the entirety of the MoE operator (Figure 3 bottom left). Instead of multiple kernel launches coordinated by the CPU, FlashMoE requires launching only one kernel, significantly reducing the involvement of the CPU. Within the fused kernel, FlashMoE implements a reactive programming model to achieve fine-grained parallelism and loosely coupled, non-blocking execution among tens of thousands of GPU threads.
MoE Implementation GPU Ops
FlashMoE (this work) 1
COMET [comet] 33
Megatron-LM CUTLASS [megatron, 10.1145/3458817.3476209] 85
Megatron-LM TE [megatron, 10.1145/3458817.3476209] 261
Megatron-LM + DeepEP [deepep] 432
DeepSpeedMoE [pmlr-v162-rajbhandari22a] 550
Table 1: We report number of GPU operations launched by MoE implementations by profiling with Nsight Systems [nsight-metrics]. We count operations in a single MoE layer (Gate $\rightarrow$ Dispatch $\rightarrow$ Expert $\rightarrow$ Combine) on 2 A100 GPUs, where each GPU has 32 experts. FlashMoE is the first to fully fuse the distributed MoE layer into a single GPU kernel.
@@
In-kernel Block scheduling and Tile parallelism. FlashMoE implements tile-level parallelism, meaning it partitions input token matrices into smaller, independent units called tiles, which are processed by blocks but managed (scheduled and constructed) by warps. We specialize every thread block, except one, as processors to perform compute. In addition, we designate a dedicated Operating System (OS) block (4 warps) to perform administrative tasks of (1) scheduling computational work to processors (scheduler), and (2) decoding computational tasks from messages received from other GPUs (subscriber). This design allows FlashMoE to dynamically assign tasks to GPU blocks based on readiness, ensuring that no GPU SM remains idle throughout the lifetime of the MoE operator. FlashMoE selects tile dimensions to maximize GPU arithmetic intensity while benefitting from a high-degree of parallelism.
@@
Asynchronous and payload-efficient communication. By redesigning the MoE operator from the ground up, FlashMoE resolves fundamental inefficiencies inherent in the conventional MoE execution pipeline. One notable inefficiency is token padding during communication. To simplify programming complexity and due to symmetry constraints of collective communication APIs, existing implementations have to zero-pad token payloads to match predefined buffer sizes. This occurs when tokens are asymmetrically routed to experts, resulting in GPUs receiving much less than the expected capacity. However, these null payloads waste communication bandwidth, bloat data transfer latency and may lead to unnecessary computations on null matrices in some implementations. FlashMoE introduces payload-efficient communication by sending non-padded tokens only to GPUs with actively selected experts, conserving both communication and computational resources.
@@
Technical challenges. Realizing the single-kernel design of FlashMoE required solving several technical challenges to achieve high performance: (1) lightweight computational dependency management; (2) navigating optimal SM occupancy configurations; (3) implementing in-device BLAS operations; (4) minimizing inter- and intra-device synchronization overheads; (5) implementing transfer-awareness by leveraging DMA over Unified Virtual Addressing (UVA) when available. In addressing these challenges, FlashMoE’s design presents a radical departure from traditional synchronous AlltoAll collectives, where GPUs exhibit significant idle time during layer execution. For device-initiated communication, FlashMoE uses NVSHMEM [nvshm] to establish a global address space across all GPUs for Direct Memory Access (DMA) communication. For in-device BLAS, FlashMoE develops custom high-performance GEMM operations via CUTLASS [Thakkar_CUTLASS_2023].
[…]
@@ 2 Motivation
Figure 4: 4(a) shows GPU utilization averaged across 100 MoE forward passes on 2 A100s with 300 GB/s unidirectional bandwidth, where we observe up to 90% idle time, due to kernel launch gaps and non-overlapping communication.
Synchronous Communication. AlltoAll or AllGather communication as currently used in MoE frameworks is a synchronous collective operation, whose completion requires the participation of all involved GPUs. Here, disparities in processing speeds or kernel scheduling among GPUs induce a straggler effect detrimental (Figure 13) to (1) the collective operation’s performance and (2) E2E performance, as stalled GPUs cannot proceed to downstream dependent or independent tasks until the collective terminates. We expound on this problem in §A.
Kernel Launch Overhead. We compare the kernel launch overheads between FlashMoE and existing baselines. Table 1 shows the number of kernel launches during a single forward pass: FlashMoE launches exactly one persistent kernel, while the baselines launch up to 550 short-lived kernels to perform the same computation. Figure 4 provides a visual comparison using CUDA API traces captured by NSight Systems, illustrating the difference between FlashMoE and DeepEP. DeepEP exhibits many small CUDA API calls, with frequent stalls between individual operators, leading to increased GPU idle time (Figure 4(a)). In contrast, FlashMoE maintains high GPU utilization by avoiding launch overhead and synchronization gaps—achieving 93.17% GPU utilization compared to 14% for DeepEP. See §4 for experimental results and §B for a discussion of related work.
@@ 3 Fused MoE Kernel Design
Figure 5: FlashMoE Fused Kernel
Modern distributed MoE systems suffer from two limitations: (1) frequent many-to-many (AlltoAll or AllGather) collectives on the critical path, and (2) significant overhead from repeated kernel launches. We address these in FlashMoE, a fully fused MoE operator implemented as a single persistent GPU kernel. Unlike previous approaches [comet, deepep, pmlr-v162-rajbhandari22a, megatron, MLSYS2023_5616d34c, MLSYS2024_339caf45, 10.1145/3503221.3508418, 10.1145/3588964, 10.1145/3627703.3650083, 10.1145/3710848.3710868, NEURIPS2022_67d57c32], FlashMoE is the first solution to implement a completely fused Distributed MoE kernel, eliminating kernel launch overhead entirely by requiring only a single kernel launch (see Table 1).
@@ 4     $\mathbf{Dispatch}(T_{\phi},A)$
Algorithm 1  FlashMoE Distributed MoE Fused Kernel
@@
Figure 6: DMoE Functional Dependencies Expressed as a Chain of Actor Interactions. We denote $S_{b}$, $S_{h}$, and $P$ as the Subscriber, Scheduler and Processor actors, respectively. For any actor $a\in\{S_{b},\>S_{h},\>P\}$, $a^{i}$ identifies an actor on GPU ii. We define $D^{j}_{i}$ as the operator, where GPU $j$ dispatches packets of tiles to GPU ii. This diagram expresses task dependencies at the granularity of a tile, namely $GEMM_{0}$, $GEMM_{1}$, combine and communication produce an output tile. Notifications occur as signals propagated through shared memory (subscriber $\leftrightarrow$ scheduler) or global memory (scheduler $\leftrightarrow$ processor or inter-GPU communication). Note one-sided inter-GPU transfers (packet or single tile) are coupled with a signal to notify $S_{b}^{j}$ on the receiving GPU $j$ of the message’s delivery.
@@
Actor-based model. The design of FlashMoE is based on the actor model of concurrent computation [agha:85, 10.5555/1624775.1624804, Greif:75]. We implement this model by specializing GPU thread blocks and warps into three distinct actor roles: (1) ProcessorF.1), (2) SubscriberF.3), and (3) SchedulerF.2). The Processor performs compute (GEMMs and element-wise operations) and tile communication. We use CUTLASS [Thakkar_CUTLASS_2023] as the underlying infrastructure for high-performance BLAS routines and NVSHMEM for kernel-initiated communication [nvshm]. The Subscriber and Scheduler perform administrative functions. Specifically, the Scheduler assigns computational tasks to available thread blocks. Our key innovation is making the Scheduler both multithreaded, enabling high scheduling throughput, and work-conserving, ensuring consistently high GPU SM utilization. On the other hand, the Subscriber decodes tile packets from peer GPUs to task descriptors (§3.1). Of the $N$ thread blocks on a GPU, we specialize N1N-1 to adopt the Processor role. We specialize the last block as the Operating System (OS). Within this block, we specialize three warps for the Subscriber role and one warp for the Scheduler role. This split of thread blocks across actors is intentional: our goal is to use few resources for administrative tasks while reserving bulk of the resources for performing MoE computation tasks. Figure 5 summarizes the FlashMoE architecture and its constituent actors, while Algorithm 1 gives a very close translation of the system in code. Note that $A\in\mathbb{R}^{S\times H}$ is the input token matrix; $O\in\mathbb{R}^{S\times H}$ the output matrix; and $X\in\mathbb{R}^{E\times H\times D}$ is a 3-D tensor of expert weights, where $E$ denotes the number of local experts for the executing GPU, $H$ is the embedding dimension, $D$ is the FFN intermediate dimension and $S$ is the sequence length. $T_{\phi}\in\left(\mathbb{N}\times\mathbb{R}\right)^{E\times C}$ is a routing table data structure, where $T_{\phi}\left(e,c\right)=(i,w)$ indicates that token $i$ at slot $c$ dispatches to expert $e$. $w$ is the combine weight (Equation 2) and $C$ is expert capacity. The tuple structure of $T_{\phi}$ is an implementation detail. $G_{\phi}\in\mathbb{R}^{S\times E}$ captures the affinity scores produced by the gate (Equation 3). Inter-actor interactions in FlashMoE. FlashMoE decomposes MoE computation and communication at the granularity of a tile, a statically sized partition of a tensor, to achieve parallel execution and efficient overlap of tasks. Each tile maps to a discrete unit of work encapsulated by a task descriptor. The Subscriber decodes these task descriptors from the remote tile packets it receives. Concurrently, the Scheduler receives notifications about available tasks and dispatches them for execution to Processor actors that perform computations defined by these tasks, namely the feed-forward network (FFN) and expert-combine operations. Figure 6 show the chain of actor interactions, demonstrating how FlashMoE enforces DMoE functional dependencies.
@@
Determining tile dimensions in FlashMoE. Selecting appropriate tile dimensions in FlashMoE is crucial to ensure efficient GPU utilization. An undersized tile underutilizes the GPU, while excessively large tiles create register pressure, causing performance-degrading register spills to local memory. After careful parameter sweeps, we choose tile dimensions of (128, 64). Our key insights are: increasing tile width significantly raises the register usage per thread, potentially triggering costly spills; increasing tile height without adjusting thread count increases workload per thread, harming performance. Raising the thread count per block beyond our fixed value of 128 threads reduces the number of concurrent blocks, negatively affecting SM occupancy. Larger thread-block sizes also increase overhead from intra-block synchronization (__syncthreads() barriers), further degrading performance. Thus, our chosen tile dimensions balance register usage, shared-memory constraints, and GPU occupancy to deliver optimal performance.